the intensity of sound wave gets reduced by 20% on passing through a slab .the reduction in intensity on passing through wo consecutive slabs is

Grade:9

1 Answers

Yash Mandil
9 years ago
instead of { wo } it must be two
let the intensity of sound be 100
after passing through a slab,
percentage reduced = 20%
amount of intensity reduced = 20% of 100 = 20
remaining intensity = 100 – 20 = 80
after passing through another slab,
intensity of sound = 80
percentage reduction = 20%
amount of intensity reduced = 20% of 80 = 16
remaining intensity = 80 – 16 = 64
total amount of intensity reduced = 100 – 64 = 36
reduction percentage = 36/100 * 100 = 36%